OpenStudy (anonymous):

The area of a rectangle is 4,320 ft2. The ratio of the length to the width is 6:5. Find the length of the rectangle. A. 12 ft B. 60 ft C. 72 ft D. 132 ft

OpenStudy (anonymous):

Okay, we need the formula for calculating the Area of a triangle. Any idea?

OpenStudy (anonymous):

i dont got it

OpenStudy (anonymous):

Well, you need to make an effort, look for it.

OpenStudy (anonymous):

let me give you another question forget about this one

OpenStudy (anonymous):

deal with this one first Robert

OpenStudy (anonymous):

Are you on a test?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

If AB:BC is 3:11, solve for x. A. 6 B. 9 C. 22 D. 28

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i dont know how to put it together

OpenStudy (anonymous):

Sorry, that's wrong. It's this\[Area \Delta = \frac{ 1 }{ 2 } bh\]

OpenStudy (anonymous):

1/2 4,320*6:5 like that

OpenStudy (anonymous):

I'm wrong again. Question says rectangle. Lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So it's Area of Rect. = Length x width

OpenStudy (anonymous):

so its 4,320=6.5*6.5

OpenStudy (anonymous):

Now what we're given the ration of the length to width to be 6:5. That means \[\frac{ l }{ b } = \frac{ 6 }{ 5 }\]We get value of l by multiplying both sides by b \[l = \frac{ 6}{ 5 }b\]

OpenStudy (anonymous):

or going right along the path we chose, calling breadth, width \[l = \frac{ 6}{ 5 }w\]

OpenStudy (anonymous):

So that value will be subst in original eqn \[Area = l \times b\] Area = 4320ft^2, we know what l is in terms of b Substituting...\[4320 = \frac{ 6}{ 5 }w \times w\] \[21600 = 6w^2\]Div both sides by 6\[3600 = w^2\]Take square roots of both sides of eqn\[w = 60ft\] So, now we've got the width. Going back to the ealier eqn we derived \[l = \frac{ 6}{ 5 }w\]Substituting w to get l\[l = \frac{ 6}{ 5 } \times 60 = 72ft\]

OpenStudy (anonymous):

Let me know if you need any clarifications.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw :)